OpenStudy (anonymous):

The atomic mass of potassium is 39.1. What is the mass of 6.02 × 1023 atoms of potassium? A. 39.1 mg B. 39.1 g C. 39.1 kg D. 6.02 × 39.1 mg E. 6.02 × 39.1 g

OpenStudy (anonymous):

Atomic mass = Mass of 1 atom = 39.1 amu Mass of 6.02 × 1023 atoms = 39.1 x 6.02 x 10^23 amu Btw, 6.02 x 10^23 atoms = 1 mol Number of mol = Mass (in gram) / Atomic Mass 1 = Mass (in gram) / 39.1 Mass (ingram) = ?

OpenStudy (anonymous):

the answer is b